The distance between the two slits in a young double slit experiment is d and the distance of the screen from the plane of the slit is b,P is a point on the screen directly in front of one of the slits the path difference between the waves arriving between at P from the two slits is

Grade:12

1 Answers

Kushagra
3 years ago
Let, y=vertical distance on the screen
      
   y/b=(path difference)/d
   y=d/2 (as the P is in front of one of the slits)
   so,
       path difference=d2/2b